OpenStudy (anonymous):

Divide and Simplify: Please help!!! √(250x^16) / √(2x)

OpenStudy (anonymous):

start with \[\frac{250x^{16}}{2x}=150x^{15}\]

OpenStudy (anonymous):

then \[\sqrt{150x^{15}}=5x^7\sqrt{6x}\]

OpenStudy (anonymous):

how did you get 5x^7 sq 6x?

OpenStudy (anonymous):

?